OpenStudy (anonymous):

is there a value for h that makes it possible for the equation square root x+h+5= 0 to have any real number solutions

OpenStudy (anonymous):

This equation? \[ \sqrt{x+h+5} = 0 \]

OpenStudy (anonymous):

Yes. for (x+h+5)^1/2 to be real x+h+5 has to be postie. Otherwise you're dealing with imaginary numbers. Therefore x+h+5>=0 x+h>-5